# Towards Unpolarized GPDs from Pseudo-Distributions

### v1

@@ 5.1 Local matrix elements
The next step of our analysis consists in forming bins of data in $t$, as indicated by the grey regions in Fig. 7. This is beneficial for several reasons. First, by averaging over data with various $(\vec{p}_{f},\vec{p}_{i})$ in a single bin, we hope to mitigate some lattice systematic uncertainty. For instance, we have many cases of symmetry $\xi\longleftrightarrow-\xi$ in our data, but also different magnitudes of $(\vec{p}_{f},\vec{p}_{i})$ in the same bin producing different excited state behavior. Furthermore, the binning results in a decrease in the number of degrees of freedom for the final fit of the $t$-dependence, meaning that the empirical covariance matrix is a better estimate of the true covariance of the data. Indeed, to estimate the covariance matrix, one should typically have many more measurements than random variables. Here, with 186 kinematic values for 358 gauge configurations, the stability of the empirical covariance matrix is reduced. Trimming down the 186 kinematic values to fifteen bins improves the confidence in this estimation. Finally, the binning gives us access to a significant number of values of $(\nu,\xi\nu)$ in each bin when $z\neq 0$, which will allow for the extraction of generalized form factors in the next section.
@@
Next, we displace the points in each bin towards the center of their bin using the $t$-slope indicated by the dipole fit. With this method, we expect to correct the intrabin $t$-dependence of order $\mathcal{O}(t-t_{\textrm{center}})$, where $t_{\textrm{center}}$ is the value at the center of the bin to which each point belongs. In effect, the modification is generally very small as $|t-t_{\textrm{center}}|$ is typically below 0.02 GeV2. A few reach 0.04 GeV2, but only for larger values of $t$, where the $t$ dependence is reduced. The only noticeable effect of intrabin $t$-correction arises in the first bin, where the slope in $t$ is pronounced and the data very numerous, and we consider this therefore as a negligible effect in the overall uncertainty budget. For this reason, we will not quote an uncertainty in $t$ in our results.
@@ 5.2 Elastic form factors from non-local data
For instance, despite the clear trend in $z$ that appears visually, one could wish to fit our dataset by a constant. Intuitively, one would expect the constant to fall somewhere in the middle of the plot, around $F_{1}\approx 0.31$. Such would indeed be the case if our points were uncorrelated. But the best fit by a constant is in fact $F_{1}\approx 0.28$, strangely close to the result at $z=0$, and in good agreement with the projection to $z=0$ of the quadratic fit by $\alpha+\beta(z/a)^{2}$. Excluding the local data at $z=0$ produces exactly the same fit results, either for the constant or quadratic fit. However, now the constant fit is incompatible with every single point of the fitted dataset. The remarkable agreement of the constant fit with the quadratic one at $z=0$ is quite surprising and happens in a general fashion in other kinematics as well. Although we will use the quadratic fit for the non-local elastic form factor, it is interesting to try to understand why a value that is significantly below every single data point can be the best constant fit of the dataset. This illustrates how such plots can be quite misleading if the very high degree of correlation is not taken into account.
@@
With this in mind, we can extract the elastic form factor using only non-local matrix elements, and performing either constant or quadratic fits over $z\in\{a,...,6a\}$. When using kinematics that are not the special $p_{f,3}=p_{i,3}=0$, since $\nu\neq 0$ and $\xi\neq 0$, the matrix element contains more than just the elastic form factor. But when the higher-order moments are fitted jointly, as we will describe in the next section, the entire kinematic dataset can be used to constrain the elastic form factor using the non-local data. The result is shown in Fig. 11. In light grey, the drifting value of $\mathcal{A}_{1,4}$ with increasing $z$ is depicted, while its value at $z=0$ is represented by the red dots. The fit of the grey dots by a quadratic is depicted as the blue dots, barely distinguishable from the $z=0$ data, or the result of a constant fit. The final fit by a $z$ expansion of either the local (red) or non-local (blue) elastic form factor extraction is completely compatible, although the $\chi^{2}$ per d.o.f. are typically significantly smaller when using the non-local extraction. We report the results of the non-local extraction of elastic form factors in Table 2 as well, and the dipole result serves as the standard result reported at the beginning of this document in Fig. 2. One will notice that the pattern of spurious $z^{2}$ dependence is clearly different between the $F_{1}$ and $F_{2}$ elastic form factors.
Figure 11: In grey, the series of 6 points $\mathcal{A}_{1,4}(\nu=0,t,z^{2})$ for $z\in\{a,...,6a\}$, in blue a quadratic (correlated) fit of the grey points, in red the data at $z=0$. The grey points are shifted gradually in $t$ to allow us to observe their trend. For instance, at large $t$, there is no spurious $z^{2}$ drift in $F_{2}$ on the right panel. The fit by a $z$-expansion is very similar whether one uses the local or non-local extraction. We merge the two cuts in lattice time separation.
@@ 6 Generalized Form Factors
up to power corrections and lattice spacing systematic uncertainties. A similar relation is obtained by substituting the appropriate combination to form the $E$ GPD $\mathcal{A}_{4}+\nu\mathcal{A}_{6}-2\xi\nu\mathcal{A}_{7}$ and its $B_{n,k}$ generalized form factors. We fit this relation at a fixed value of $z$ in each bin, with an intrabin $t$ correction proceeding in the same fashion as what we have described for the elastic form factors. The result using all points in the bin $t\approx-0.33$ GeV2, $z=6a$, and the cut $t_{s}=3$ is displayed in Fig. 12. The $\chi^{2}$ per d.o.f. of this fit of 28 points with 3 parameters is close to 2 both for the real and imaginary parts. Note that although it may seem that the special kinematic point at $\nu=0$ corresponding to $\vec{p}_{f}=(0,0,0)$ and $\vec{p}_{i}=(1,1,0)$ is dragging the value of $F_{1}(t)$ upwards, this is not the case. Excluding this point from the fit would give $F_{1}(t,z=6a)=0.687\pm 0.008$, whereas with it included, we find the very similar $F_{1}(t,z=6a)=0.689\pm 0.007$. As we have already discussed, the data favor a value of $F_{1}(t)$ at $z=6a$ that is noticeably larger than the data at $z=0$ prefers, namely $0.652\pm 0.007$, that is $5\sigma$ below. One will also notice that data at larger $\xi$ exists at smaller values of $\nu$ since part of the average $P_{3}$ momentum is sacrificed to create the momentum transfer. The coverage in $(\xi,\nu)$ available for our study is depicted in Fig. 13, with the exception of a few points where $p_{f,3}=-p_{i,3}\neq 0$, corresponding to $\nu=0$ and $\xi$ infinite (but $\nu\xi\neq 0$). As one can see from Eqs. (78) and (79), those kinematics probe exclusively the elastic form factor and the skewness-dependent generalized form factors, so they are particularly useful.
@@ 6.1 Matching
The full leading-logarithm calculation exhibits the smallest scale fixing uncertainties of the three matching procedures, the increasingly large $\alpha_{s}$ balancing one another in the kernel. The procedure produces relatively constant results up to $z=0.3$ fm, after which it is no longer constant in this bin. In principle, we do not have expectations that a perturbative matching explains the $z^{2}$ dependence of the data when $-1/z^{2}\ll 1$ GeV2. Therefore, this situation is not particularly concerning. In fact, when applied to the $E$ GPD in the same bin, the full leading-logarithmic matching still has the smallest scale fixing uncertainties, and overall the best behavior as seen in Fig. 15. At this stage, considering that different behaviors are observed for various bins, no strong conclusion can be drawn, especially in view of the spurious $z$ dependence observed in the elastic form factors.
@@
Figure 16: The moments of the GPDs $H^{u-d}$ (left) and $E^{u-d}$ (right) at $t=-0.33$ GeV2 with leading-logarithmic matching to $\mu=2$ GeV as a function of $z^{2}$ (cut $t_{s}=4$). In the upper row, we show the skewness-independent moments, and in the lower row the skewness-dependent moments.
@@ 6.2 $t$-dependent generalized form factors
The results of the fits of the $t$-dependence of moments of GPDs are presented in Fig. 17 for the generalized form factors $A_{n,0}$ and $B_{n,0}$ that are independent of $\xi$, and in Fig. 18 for the $A_{n,2}$ and $B_{n,2}$ that are dependent on $\xi$. Since beyond the elastic form factor, the trend in $z$ of the data is not discernible in a consistent fashion, we use a constant fit on $z\in\{a,...,6a\}$ after having matched the data to $\overline{MS}$ at 2 GeV using the leading-logarithmic matching derived in section 2.3. The colored bands result from the statistical uncertainty of the fit on data with a cut in lattice time separation of three. The uncolored band results from the union of the result with a cut of three and a cut of four. It appears that excited state uncertainty increases as we head towards larger moments, unsurprisingly as the signal in the data weakens. In the specific case of $A_{4,2}$ and $B_{4,2}$, the dipole fit of the data with a cut of four is extremely unstable, as some data (especially in the cut of four) is compatible with 0 or negative. As a non-linear fit is unable to cross zero, the dipole fit has major issues adapting to these data. We use there therefore the $z$-expansion fit, which on the other hand does not suffer from any particular difficulty, to increase the stability of the dipole. The obtained results for the dipole fit in terms of value at $t=0$ and mass are reported in Fig. 2.
Figure 17: The skewness-independent generalized form factors from $\int\mathrm{d}x\,xF$ to $\int\mathrm{d}x\,x^{3}F$, as a function of $t$. The data points are obtained for a time separation cut of three in the analysis of correlation functions. The fit of those data points, either by a dipole or a $z$ expansion, is represented by the uniformly colored areas. When merging this uncertainty with the one using a time separation cut of four, one obtains the wider uncertainties, which contain information about excited state contamination.
Figure 18: The skewness-dependent generalized form factors. Same caption as Fig. 17.
@@ 6.3 Radial distributions
We use in the following the dipole fits, as the $t$-dependence of the $z$ expansion is too unconstrained at large $t$ to give rise to sensible radial distributions. The extraction presented here is therefore strongly model dependent. Even then, the $t$-dependence of $A_{4,0}$ and $B_{4,0}$ is very poorly determined and therefore we only present results for $n\in\{0,1,2\}$.
@@
The unpolarized moments (85) are spherically symmetric in $|\vec{b}_{\perp}|$. We represent them in Fig.  19. As can already be inferred by the increasing value of the fitted dipole mass (see the summary table of Fig.  2), when the order of the skewness-independent moments increases, the radial extent of the distribution shrinks. This can be largely understood from a kinematic point of view, as higher-order Mellin moments are increasingly dominated by the large $x$ domain of the distribution. The center of the proton with respect to which our radial coordinate system is defined is the barycenter of the partonic longitudinal momentum. When the longitudinal momentum fraction of the active parton $x$ approaches one, the active parton is by necessity close to the center of the proton, so the impact parameter distribution at large $x$ is very narrow.
@@ 7 Conclusions
With the current statistical precision, control over excited states is critical for an accurate analysis. We have compared results from two different methods for extracting matrix elements and from different cuts on the data. This step is critical for creating realistic error estimates for the matrix elements. The data with different cuts in Euclidean time can lead to significant deviations in the final results, as was seen in the extraction of generalized form factors. In future studies, we will use the distillation to implement distinct nucleon operators for each momentum. This allows for a GEVP analysis to explicitly remove contributions from excited states. The excited state contamination will only become more difficult to control in the chiral limit, so this full control of excited state contamination will be crucial for physical extrapolations.

### v2

@@ 5.1 Local matrix elements
The next step of our analysis consists in forming bins of data in $t$, as indicated by the grey regions in Fig. 7. This is beneficial for several reasons. First, by averaging over data with various $(\vec{p}_{f},\vec{p}_{i})$ in a single bin, we hope to mitigate some lattice systematic uncertainty. For instance, we have many cases of symmetry $\xi\longleftrightarrow-\xi$ in our data, but also different magnitudes of $(\vec{p}_{f},\vec{p}_{i})$ in the same bin producing different excited state behavior. Furthermore, the binning results in a decrease in the number of degrees of freedom for the final fit of the $t$-dependence, meaning that the empirical covariance matrix is a better estimate of the true covariance of the data. Indeed, to estimate the covariance matrix, one should typically have many more measurements than random variables. Here, with 186 kinematic values for 348 gauge configurations, the stability of the empirical covariance matrix is reduced. Trimming down the 186 kinematic values to fifteen bins improves the confidence in this estimation. Finally, the binning gives us access to a significant number of values of $(\nu,\xi\nu)$ in each bin when $z\neq 0$, which will allow for the extraction of generalized form factors in the next section.
@@
8 Next, we displace the points in each bin towards the center of their bin using the $t$-slope indicated by the dipole fit. With this method, we expect to correct the intrabin $t$-dependence of order $\mathcal{O}(t-t_{\textrm{center}})$, where $t_{\textrm{center}}$ is the value at the center of the bin to which each point belongs. In effect, the modification is generally very small as $|t-t_{\textrm{center}}|$ is typically below 0.02 GeV2. A few reach 0.04 GeV2, but only for larger values of $t$, where the $t$ dependence is reduced. The only noticeable effect of intrabin $t$-correction arises in the first bin, where the slope in $t$ is pronounced and the data very numerous, and we consider this therefore as a negligible effect in the overall uncertainty budget. For this reason, we will not quote an uncertainty in $t$ in our results.
@@ 5.2 Elastic form factors from non-local data
For pedagogical purpose on the counter-intuitive effect of very high correlation, despite the clear trend in $z$ that appears visually, one could wish to fit our dataset by a constant. Intuitively, one would expect the constant to fall somewhere in the middle of the plot, around $F_{1}\approx 0.31$. Such would indeed be the case if our points were uncorrelated. But the best fit by a constant is in fact $F_{1}\approx 0.28$, strangely close to the result at $z=0$, and in good agreement with the projection to $z=0$ of the quadratic fit by $\alpha+\beta(z/a)^{2}$. Excluding the local data at $z=0$ produces exactly the same fit results, either for the constant or quadratic fit. However, now the constant fit is incompatible with every single point of the fitted dataset. The remarkable agreement of the constant fit with the quadratic one at $z=0$ is quite surprising and happens in a general fashion in other kinematics as well. Although we will use the quadratic fit for the non-local elastic form factor, it is interesting to try to understand why a value that is significantly below every single data point can be the best constant fit of the dataset. This illustrates how such plots can be quite misleading if the very high degree of correlation is not taken into account.
@@
With this in mind, we can extract the elastic form factor using only non-local matrix elements, and performing quadratic fits over $z\in\{a,...,6a\}$. When using kinematics that are not the special $p_{f,3}=p_{i,3}=0$, since $\nu\neq 0$ and $\xi\neq 0$, the matrix element contains more than just the elastic form factor. But when the higher-order moments are fitted jointly, as we will describe in the next section, the entire kinematic dataset can be used to constrain the elastic form factor using the non-local data. The result is shown in Fig. 11. In light grey, the drifting value of $\mathcal{A}_{1,4}$ with increasing $z$ is depicted, while its value at $z=0$ is represented by the red dots. The fit of the grey dots by a quadratic is depicted as the blue dots, barely distinguishable from the $z=0$ data, or the result of a constant fit. The final fit by a $z$ expansion of either the local (red) or non-local (blue) elastic form factor extraction is completely compatible, although the $\chi^{2}$ per d.o.f. are typically significantly smaller when using the non-local extraction. We report the results of the non-local extraction of elastic form factors in Table 2 as well, and the dipole result serves as the standard result reported at the beginning of this document in Fig. 2. One will notice that the pattern of $z^{2}$ dependence is clearly different between the $F_{1}$ and $F_{2}$ elastic form factors.
Figure 11: In grey, the series of 6 points $\mathcal{A}_{1,4}(\nu=0,t,z^{2})$ for $z\in\{a,...,6a\}$, in blue a quadratic (correlated) fit of the grey points, in red the data at $z=0$. The grey points are shifted gradually in $t$ to allow us to observe their trend. For instance, at large $t$, there is no $z^{2}$ drift in $F_{2}$ on the right panel. The fit by a $z$-expansion (79) is very similar whether one uses the local or non-local extraction. We merge the two cuts in lattice time separation.
@@ 6 Generalized Form Factors
up to power corrections and lattice spacing systematic uncertainties. A similar relation is obtained by substituting the appropriate combination to form the $E$ GPD $\mathcal{A}_{4}+\nu\mathcal{A}_{6}-2\xi\nu\mathcal{A}_{7}$ and its $B_{n,k}$ generalized form factors. The question of the $D$-term extracted from $\mathcal{A}_{5}$ is discussed in section 6.4. We fit these relation at a fixed value of $z$ in each bin, with an intrabin $t$ correction proceeding in the same fashion as what we have described for the elastic form factors. The result using all points in the bin $t\approx-0.33$ GeV2, $z=6a$, and the cut $t_{s}=3$ is displayed in Fig. 12.
@@
The $\chi^{2}$ per d.o.f. of this fit of 28 points with 3 parameters is close to 2 both for the real and imaginary parts. Note that although it may seem that the special kinematic point at $\nu=0$ corresponding to $\vec{p}_{f}=(0,0,0)$ and $\vec{p}_{i}=(1,1,0)$ is dragging the value of $F_{1}(t)$ upwards, this is not the case. Excluding this point from the fit would give $F_{1}(t,z=6a)=0.687\pm 0.008$, whereas with it included, we find the very similar $F_{1}(t,z=6a)=0.689\pm 0.007$. As we have already discussed, the data favor a value of $F_{1}(t)$ at $z=6a$ that is noticeably larger than the data at $z=0$ prefers, namely $0.652\pm 0.007$, that is $5\sigma$ below. One will also notice that data at larger $\xi$ exists at smaller values of $\nu$ since part of the average $P_{3}$ momentum is sacrificed to create the momentum transfer. The coverage in $(\xi,\nu)$ available for our study is depicted in Fig. 13, with the exception of a few points where $p_{f,3}=-p_{i,3}\neq 0$, corresponding to $\nu=0$ and $\xi$ infinite (but $\nu\xi\neq 0$). As one can see from Eqs. (85) and (86), those kinematics probe exclusively the elastic form factor and the skewness-dependent generalized form factors, so they are particularly useful.
@@ 6.1 Matching
The full leading-logarithm calculation exhibits the smallest scale fixing uncertainties of the three matching procedures. Although $\alpha_{s}(\lambda/z^{2})$ increases considerably at large $z$, since $\gamma_{1}$ is negative, the ratio $(\alpha_{s}(\lambda/z^{2})/\alpha_{s}(\mu^{2}))^{\gamma_{1}C_{F}/(2\pi\beta_{0})}$ decreases. On the other hand, the final factor of Eq. (89) increases at large $z$, and both variations happen in quite similar proportions. The procedure produces relatively constant results up to $z=0.3$ fm, after which it is no longer constant in this bin. In principle, we do not have expectations that a perturbative matching explains the $z^{2}$ dependence of the data when $-1/z^{2}\ll 1$ GeV2. Therefore, this situation is not particularly concerning. In fact, when applied to the $E$ GPD in the same bin, the full leading-logarithmic matching still has the smallest scale fixing uncertainties, and overall the best behavior as seen in Fig. 15. At this stage, considering that different behaviors are observed for various bins, no strong conclusion can be drawn, especially in view of the $z^{2}$ dependence observed in the elastic form factors.
@@
Figure 16: The moments of the GPDs $H^{u-d}$ (left) and $E^{u-d}$ (right) at $t=-0.33$ GeV2 with leading-logarithmic matching to $\mu=2$ GeV as a function of $z/a$ (cut $t_{s}=4$). In the upper row, we show the skewness-independent moments, and in the lower row the skewness-dependent moments.
@@ 6.2 $t$-dependent generalized form factors
The results of the fits of the $t$-dependence of moments of GPDs are presented in Figs. 17 to 19 for the generalized form factors $A_{n,0}$ and $B_{n,0}$ that are independent of $\xi$, and in Fig. 20 for the $A_{n,2}$ and $B_{n,2}$ that are dependent on $\xi$. Since beyond the elastic form factor, the trend in $z$ of the data is not discernible in a consistent fashion, we use a constant fit on $z\in\{a,...,6a\}$ after having matched the data to $\overline{MS}$ at 2 GeV using the leading-logarithmic matching derived in section 2.3. We demonstrate in Appendix C that the correlated constant fit gives very robust results on a diverse subset of separations, for instance using only $z\in\{a,...,4a\}$. The colored bands are 68% confidence intervals obtained by dropping the lower and upper 16% of the best-fit results on the data with a cut in lattice time separation of three. This makes practically no difference compared to a simple $1\sigma$ uncertainty band, except for the most poorly determined moments $B_{4,0}$ and $A_{4,2}$, $B_{4,2}$ where the non-linearity of the dipole fit becomes apparent and the distributions of best fits are non-Gaussian. The uncolored band results from the union of the result with a cut of three and a cut of four. It appears that excited state uncertainty increases as we head towards larger moments, unsurprisingly as the signal in the data weakens. The obtained results for the dipole fit in terms of value at $t=0$ and mass are reported in Fig. 2.
Figure 17: The generalized form factors $A_{2,0}$ (left) and $B_{2,0}$ (right) as a function of $t$. The data points are obtained for a time separation cut of three in the analysis of correlation functions. The fit of those data points, either by a dipole or a $z$ expansion, is represented by the uniformly colored areas. When merging this uncertainty with the one using a time separation cut of four, one obtains the wider uncertainties, which contain information about excited state contamination.
Figure 18: The generalized form factors $A_{3,0}$ (left) and $B_{3,0}$ (right) (see caption of Fig. 17)
Figure 19: The generalized form factors $A_{4,0}$ (left) and $B_{4,0}$ (right) (see caption of Fig. 17)
Figure 20: The skewness-dependent generalized form factors $A_{3,2}$ (top left), $B_{3,2}$ (top right), $A_{4,2}$ (bottom left) and $B_{4,2}$ (bottom right) (see caption of Fig. 17).
@@ 6.3 Radial distributions
In the following we use the dipole fits as the $t$-dependence of the $z$ expansion is too unconstrained at large $t$ to give rise to sensible radial distributions. The extraction presented here is therefore strongly model-dependent. Even then, the $t$-dependence of $A_{4,0}$ and $B_{4,0}$ is very poorly determined and therefore we only present results for $n\in\{0,1,2\}$.
@@
The unpolarized moments (92) are spherically symmetric in $|\vec{b}_{\perp}|$. We represent them in Fig.  21. As can already be inferred by the increasing value of the fitted dipole mass (see the summary table of Fig.  2), when the order of the skewness-independent moments increases, the radial extent of the distribution shrinks. This can be largely understood from a kinematic point of view, as higher-order Mellin moments are increasingly dominated by the large $x$ domain of the distribution. The center of the proton with respect to which our radial coordinate system is defined is the barycenter of the partonic longitudinal momentum. When the longitudinal momentum fraction of the active parton $x$ approaches one, the active parton is by necessity close to the center of the proton, so the impact parameter distribution at large $x$ is very narrow.
@@
6.4 $D$-term
For the isovector $D$-term, Eq. (45) invites us to fit:
$\textrm{Im}\,\mathcal{A}_{5}(\nu,\xi,t,z^{2})=\xi\nu C_{2}(t,z^{2})+\mathcal{O}(\nu^{3})+\mathcal{O}(\Lambda_{\rm QCD}^{2}z^{2},tz^{2})\,.$ (96)
As we have noticed before, there is a trade-off between the value of $\xi$ and $\nu$ as larger $\xi$ requires sacrificing some of the $P_{3}$ momentum to introduce a momentum transfer. As a result, the range in $\nu\xi$ available to extract the $D$-term does not exceed 1.2 when $z=6a$, and we can only achieve sensitivity to the $C_{2}$ moment of the $D$-term. The situation in one of the bins in $t$ is depicted in Fig. 23. Because of the absence of any discernable trend in $t$, we do not attempt an intrabin $t$ correction in the fashion we used for the rest of the GPD.
[…]
In general, $\mathcal{A}_{5}$ is compatible with 0 and small in magnitude. Some kinematics $(\vec{p}_{f},\vec{p}_{i})$ show noticeably non-zero values in the real part, a problem we have already identified when discussing the contribution of $\mathcal{A}_{5}$ in the local matrix element. Therefore, there is a clear sign of potential lattice discretization errors, which we have no means of subtracting from the data within this analysis. We also note that Ref. Radyushkin (2024) highlights the potential presence of “kinematic” higher-twists in $\mathcal{A}_{5}$. When using the full statistics available in some bins, it is possible to fit the imaginary part with the functional form of Eq. (96) and obtain an extraction for $C_{2}$ only a few standard deviations away from 0. The imaginary part of $\mathcal{A}_{5}(\nu,\xi)$ is compatible with being an odd function of the product $\nu\xi$, which mirrors the fact that the $D$-term is an odd function of $x/\xi$. The matching of $C_{2}$ then proceeds exactly similarly to the one of $A_{2,0}$ and $B_{2,0}$ (45).
The final result on $C_{2}$ is depicted in Fig. 24 matched to 2 GeV using the data $z=a$ up to $6a$. The dipole and the $z$-expansion fit produce very different $t$-dependence and extrapolation in the limit where $t\rightarrow 0$, due to the weakness of the constraint imposed by the data. The small magnitude of $C_{2}$ is consistent with the results of other lattice calculations from local matrix elements, e.g. Hagler et al. (2008); Bali et al. (2019); Alexandrou et al. (2020a). In the large $N_{c}$ approximation, one can note that the $u-d$ contribution to the $D$-term is suppressed by a factor $1/N_{c}$ compared to the $u+d$ flavor combination Goeke et al. (2001).
[…]
@@ 7 Conclusions
With the current statistical precision, control of excited state contamination is critical for an accurate analysis. We have compared results from two different methods for extracting matrix elements and from different cuts on the data. This step is critical for creating realistic error estimates for the matrix elements. The data with different cuts in Euclidean time can lead to significant deviations in the final results, as was seen in the extraction of generalized form factors. In future studies, we will use the distillation to implement distinct nucleon operators for each momentum. This allows for a GEVP analysis to explicitly remove contributions from excited states. The excited state contamination will only become more difficult to control in the chiral limit, so this full control of excited state contamination will be crucial for physical extrapolations.
@@
The use of non-local separations in the matrix elements up to $z=0.6$ fm gives rise to challenges for the proper evaluation of the systematic uncertainty caused by the perturbative matching, as well as the contamination by higher twist contributions, which we do not have the possibility to properly treat in this study. In view of some clear signal of lattice discretization errors, which we cannot easily differentiate from higher twist contributions, we see the realization of a continuum limit of this study as an important further milestone.